Find the x- and y-intercepts apply elimination or substitution to find the coordinates of the point of intersection.
15x + 11y = -17
8x - 6y = 74

Answer by mananth(16946)

15x + 11y = -17................1
8x - 6y = 74...................2
multiply equation 1 by 6 and equation 2 by 11 and add them. to eliminate y
..
90x+66y=-102
88x-66y=814
178x=712
x= 712/178
x= 4
plug the value of x in any equation
90*4 +66y=-102
360+66y=-102
66y=-102-360
66y=-462
y=-7
The intercepts are (4, -7)